One thousand sheets of aluminum alloy produced by the same manufacturer were examined for surface flaws. Assume surface flaws occur randomly and independently at a constant rate with an average of 4 flaws per sheet. Sheets are all the same size.

a) [3 points] What is the probability that a randomly chosen sheet contains more than 1 surface flaw?

b) [3 points] When 10 sheets are picked at random, what is the probability that 6 of these sheets have more than 1 surface flaw on them?

c) [4 points] Given that a sheet has more than 1 surface flaw on it, find the probability that it has exactly 10 surface flaws.
Answer #1

it is poisson process :

flaws/sheet = 4/1

P(k flaws in x sheets) = (e^(-4*x))*((4*x)^k)/(k!)

a.

P(k>1 in 1 sheet) = 1 - P(k=0 in 1 sheet) - P(k=1 in 1 sheet)

= 1 - (e^(-4*1))*((4*1)^0)/(0!) - (e^(-4*1))*((4*1)^1)/(1!)

P(k>1 in 1 sheet) = 0.9084

b.

P(more than 1 flaw in sheet) = 0.9084

n=10, x=6, p=0.9084, 1-p = 0.0916

P(6) = 10C6 * (0.9084^6)* ((0.0916)^(4)) = 0.0083

P(6 sheets with more than 1 flaw) = 0.0083

c.

P(10 flaws and more than 1 flaw) = P(10 flaw) {as more than 1 flaw condition is fulfilled if flaws=10}

P(10 flaw) = (e^(-4*1))*((4*1)^10)/(10!) = 0.0053 = P(10 flaws and more than 1 flaw)

P(more than 1 flaw) = 0.9084 {from part a}

P(10 flaw | more than 1 flaw) = P(10 flaw and more than 1 flaw) / P(more than 1 flaw)

= 0.0053/0.9084

P(10 flaw | more than 1 flaw) = 0.0058

P.S. (please upvote if you find the answer satisfactory)